What is Spin-orbit coupling? Explain if you expect to observe it in a ground state Na atom (elec. conf. 1s2 2s2 sp6 3s1) or an excited Na atom with conf. 1s2 2s2 2p6 3s1?
Answer #1

The interaction of the spin magnetic moment with the magnetic field arising from orbital angular momentum (an electron in an orbital with l>0) is called Spin-Orbit coupling.

It is observed in the Na spectra that D-Line (yellow emission) is split into two lines D1 (589.59 nm) and D2 (588.99 nm). The reason is that the energy levels split into two except S level (l=0)

The term arising from G.S configuration is 2S l is 0. S=1/2 so according to 2S+1LJ we get 2S1/2

for the first excited state terms is 2P l=1 s=1/2 now j can have two values= 1/2 and 3/2. Now this term will split into two energy levels 2P1/2 and 2P3/2

So it is expected in excited Na atom.